HelpMePlz:

Need help and explaining with some rational polynomials: Solve 6/(x - 3) = 3/ x for x and determine if the solution is extraneous or not.

Champion:

Is it: \(\Large\frac{6}{x - 3} = \frac{3}{ x}\) (notice how parentheses are not required in the denominator, but are needed when we write the expression on a single line) In this case, we cross multiply and end up with a linear equation, so there is no question of extraneous roots: \(\Large 6x=3(x-3)\) Simplify, distribute, isolate x and solve for x.